OpenStudy (anonymous):

Find the x-intercept(s) and the coordinates of the vertex for the parabola y = x^2 - 2x - 24 . If there is more than one x-intercept, separate them with commas.

OpenStudy (anonymous):

x intercepts are -4 and 6

OpenStudy (anonymous):

factor and you will get them

OpenStudy (amistre64):

i like completeing the square for these; its quicker 24 = x^2 -2x 24+1 = x^2 -2x+1 25 = (x-1)^2 +- 5 = x-1 x = 1+5; 1-5 x = 6,-4

OpenStudy (amistre64):

well; its more effortless then trial and error I should say lol

OpenStudy (anonymous):

wait so if those r the x-intercepts whic is the vertex ?

OpenStudy (anonymous):

vertex use \[-\frac{b}{2a}\]

OpenStudy (amistre64):

vertex should be 1,-25 if i see it right

OpenStudy (anonymous):

this does not ask for vertex.

OpenStudy (amistre64):

"Find the x-intercept(s) and the coordinates of the vertex ..."

OpenStudy (anonymous):

half way between -4 and 6 is 1 yes

OpenStudy (amistre64):

those 2 part questions will bite you every time lol

OpenStudy (anonymous):

i 4 got 2 include it in the question its lokking for the x-intecept(s) and the vertex.....sorry

OpenStudy (anonymous):

and if x = 1 y = -25

OpenStudy (anonymous):

so the vertex is (1,-25) ?

OpenStudy (anonymous):

so intercepts are (-4,0) and (6,0) vertex is (1,-25)

OpenStudy (anonymous):

THANK YOU !!!!!!!!!!!!!!!!!!!!!!!